Grade 9Algebra

a+b=a/b+b/a. Find a*2+b*2.give the answer pleaseAnd as soon as possible

Katan
9 Years agoGrade 9

1 Answer

Arun
9 Years ago
This ia only valid if a= 1 and b = 1 OR a= -1 and b= -1Hence a^2 + b^2 = 1^2 + b^2= 2Note- this is the form of x+1/x this should be greater than and equal to 2 Or less than and equal to -2